OpenStudy (anonymous):

Solve equation 6 + x/2 = 3

OpenStudy (saifoo.khan):

\[\frac x2 = -3\] \[x = -3 \times 2\]

OpenStudy (anonymous):

6+x/2=3 x/2=3-6 x/2=-3 x=-3*2 x=-6